### v2

@@ Appendix B Sufficient Conditions for Assumption 2
$\leq$ $\ T!\times\left(\delta/T!\right)=\delta.$ (B.7)
Since $f^{w}$ can be chosen to make $\delta$ arbitrarily small, equation (B.7) implies that $f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)$ can be approximated arbitrarily closely by a polynomial $\overline{f}^{w}$ that is symmetric in $\left(x_{t},z_{t}\right)$ pairs for $t=1,2$. Furthermore, by the fundamental theorem of symmetric functions, for any $a$ on its support, $\overline{f}^{w}$ can be written as a polynomial function of the elementary symmetric functions $W\left(\cdot\right)$ of $\left(\left(x_{1},z_{1}\right),\left(x_{2},z_{2}\right)\right)$. We denote this function by $f_{\left.A\right|W}\left(\left.a\right|w\right)$ and obtain that $f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)$ can be approximated arbitrarily closely by $f_{\left.A\right|W}\left(\left.a\right|W\right)$. Let $\delta\rightarrow 0$ in (B.7). [^28] Note that to let $\delta\to 0$ we need the order of polynomials in the approximating function $f^{w}$ to increase. However, it does not affect the order of the elements of $W$ since by the fundamental theorem of symmetric functions, the elements of $W$ has a fixed order of $T$. The key is we have two different orders of polynomials; one is the order of the approximating functions $f^{w}$, the other is the order of the arguments of $W$. For example, one may let the order $P_{1}$ and $P_{2}$ of a polynomial function $h\left(x,y\right)=\sum_{i=1}^{P_{1}}\left(x+y\right)^{i}+\sum_{i=1}^{P_{2}}\left(xy\right)^{i}$ to increase while keeping the order of its symmetric arguments (in this example, it is $x+y$ and $xy$) fixed. See footnote 9 and 10 of Altonji and Matzkin (2005) for a more detailed discussion. Then, for any $t\in\left\{1,...,T\right\}$ and on-support $\left(x_{t},z_{t},a,w\right)$
@@ Appendix C Simulation
In this section, I examine the finite-sample performance of the series estimators via a simulation study. A discussion of the data generating process (DGP) motivated by production function applications is first provided. Then, I present the baseline results of estimating the APE $\overline{b}\coloneqq\mathbb{E}\beta_{it}$. Next, I show how the method performs for estimating the LAR $b_{t}\left(x\right)\coloneqq\mathbb{E}\left[\left.\beta_{it}\right|X_{it}=x\right]$. Finally, as robustness checks, I (i) vary the number of firms and periods, (ii) change the order of basis functions used for series estimation, and (iii) include ex-post shocks $\epsilon_{it}$ and $\upsilon_{it}$ to the DGP.
@@ C.1 DGP
$Y_{it}=k_{it}\beta_{it}^{K}+l_{it}\beta_{it}^{L}+\omega_{it}+\epsilon_{it},$
[…]
$k_{it}$ $=\dfrac{r_{it}-\left(A_{i}+\eta_{it}^{K}\right)\left(r_{it}-w_{it}\right)/10-\ln\left(\left(A_{i}+\eta_{it}^{K}\right)/10\right)-\ln\left(A_{i}+\eta_{it}^{K}/2+1\right)}{\left(A_{i}+\eta_{it}^{K}\right)/5-1},$
$l_{it}$ $=\dfrac{w_{it}-\left(A_{i}+\eta_{it}^{L}\right)\left(w_{it}-r_{it}\right)/10-\ln\left(\left(A_{i}+\eta_{it}^{L}\right)/10\right)-\ln\left(A_{i}+\eta_{it}^{L}/2+1\right)}{\left(A_{i}+\eta_{it}^{L}\right)/5-1}.$
Let $X_{it}=\left(k_{it},l_{it}\right)^{{}^{\prime}}$. It is clear that $X_{it}$ is correlated with $\beta_{it}$ in each period via $\left(A_{i},\eta_{it}^{K},\eta_{it}^{L}\right)$ in a nonseparable way. I use $N$, $T$, and $M$ to denote the total number of firms, periods, and simulations, respectively. The observed data is $\left\{X_{it,m},Y_{it,m},Z_{it,m}\right\}$ for $i=1,...,n$, $t=1,...,T$, and $m=1,...,M$, which are used to construct $\widehat{\overline{b}}_{m}$ and $\widehat{b}_{t,m}\left(x\right)$ to estimate $\overline{b}$ and $b_{t}\left(x\right)$, respectively, via the estimation procedure outlined in Section 4.1. I evaluate the performance of the estimators by their biases, root-mean squared errors (rMSE), and mean normed deviations (MND), with the explicit mathematical definitions provided in the tables below.

### v3

@@ A.2 Nonparametric Justification of the Index Exclusion Assumption
$\leq$ $\ T!\times\left(\delta/T!\right)=\delta.$ (A.75)
Since $f^{w}$ can be chosen to make $\delta$ arbitrarily small, equation (A.75) implies that $f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)$ can be approximated arbitrarily closely by a polynomial $\overline{f}^{w}$ that is symmetric in $\left(x_{t},z_{t}\right)$ pairs for $t=1,2$. Furthermore, by the fundamental theorem of symmetric functions, for any $a$ on its support, $\overline{f}^{w}$ can be written as a polynomial function of the elementary symmetric functions $W$ of $\left(\left(x_{1},z_{1}\right),\left(x_{2},z_{2}\right)\right)$. I denote this function by $f_{\left.A\right|W}\left(\left.a\right|w\right)$ and obtain that $f_{\left.A\right|X_{1},X_{2},Z_{1},Z_{2}}\left(\left.a\right|x_{1},x_{2},z_{1},z_{2}\right)$ can be approximated arbitrarily closely by $f_{\left.A\right|W}\left(\left.a\right|W\right)$. Let $\delta\rightarrow 0$ in (A.75). Note that to let $\delta\to 0$ the degree of polynomials in the approximating function $f^{w}$ needs to increase. However, it does not affect the dimension of $W$ since by the fundamental theorem of symmetric functions, the elements of $W$ have a fixed degree of $T$. The key is that there are two different degrees of polynomials: one is the degree of the approximating functions $f^{w}$, and the other is the degree of the arguments of WW. [^4] For example, one may let the degree $P_{1}$ and $P_{2}$ of a polynomial function $h\left(x,y\right)=\sum_{i=1}^{P_{1}}\left(x+y\right)^{i}+\sum_{i=1}^{P_{2}}\left(xy\right)^{i}$ to increase while keeping the degree of its symmetric arguments (in this example, $W=(x+y,xy)$) fixed. See also footnotes 9 and 10 of Altonji and Matzkin (2005) for a detailed discussion.
Then, for any $t\in\left\{1,...,T\right\}$ and on-support $\left(x_{t},z_{t},a,w\right)$
@@
A.3 Further Discussion of Identification Assumptions
When the exchangeability conditions are used to justify Assumption 2, I present a symmetry-based argument in the discussion of Assumption 4 that justifies Assumptions 2 and 4 simultaneously. I provide an example here to clarify this point. Suppose $T=2$ and $(X_{it},Z_{it})\in\mathbb{R}^{2}$. By Proposition 1, when $W_{i}$ includes averages through time of the polynomials of $\left(X_{it},Z_{it}\right)$ up to the second degree, Assumption 2 holds. Suppose $(X_{i1},X_{i2},Z_{i1},Z_{i2})=\left(x_{1},x_{2},z_{1},z_{2}\right)$ satisfies $\{V_{it}=v,W_{i}=w\}$ for $t=1,2$ where $x_{1},x_{2},z_{1},z_{2}\in\mathbb{R}^{2}$ and $x_{1}$ and $x_{2}$ are linearly independent. Then, by definition of $V_{it}$ and the symmetry of $W_{i}$ in $\left(X_{it},Z_{it}\right)$ through $t$, $(X_{i1},X_{i2},Z_{i1},Z_{i2})=\left(x_{2},x_{1},z_{2},z_{1}\right)$ must also satisfy $V_{it}=v$ and $W_{i}=w$ for $t=1,2$. Therefore, given $V_{it}=v$ and $W_{i}=w$ there are two linearly independent vectors of $x_{1}$ and $x_{2}$ that $X_{it}$ can take, satisfying Assumption 4.
However, Assumption 4 requires more residual variation in $X_{it}$ given $(V_{it},W_{i})$ and can be restrictive if nonparametric justifications are used for Assumption 2. I summarize the proposal by Altonji and Matzkin (2005) to achieve such residual variation. Specifically, Altonji and Matzkin (2005) suggest to
(i) leverage unconditional variation in the exogenous regressors that do not enter $W_{i}$, and I provide more details at the end of this section;
(ii) restrict the conditional distribution of $A_{i}$ given $W_{i}$ to depend on not all but only a subset of coordinates of $W_{i}$, which is testable by comparing the fit of $\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]$ with $W_{i}$ replaced by a proper subset of it;
(iii) assume $f_{\left.A_{i}\right|W_{i}}$ to depend on $W_{i}$ only through a linear combination of the elements of $W_{i}$, i.e., $f_{\left.A_{i}\right|W_{i}}=f_{\left.A_{i}\right|\sum_{k=1}^{d_{W}}c_{k}W_{i,k}}$ where $W_{i}=(W_{i,1},\ldots,W_{i,d_{W}})^{\prime}$ and the $c_{k}$s are unknown parameters, which implies that
$\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]=\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},\sum_{k=1}^{d_{W}}c_{k}W_{i,k}\right]$
and the $c_{k}$s can be estimated using Ichimura and Lee (1991);
(iv) impose a priori restrictions on model (2.1)–(2.2) such that $W_{i}$ enters $\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]$ in a parametric way (e.g.,
$\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right]=\sum_{k=1}^{d_{W}}c_{k}W_{i,k}+\sum_{l=1}^{d_{V}}d_{l}V_{it,l}$
when $d_{X}=1$); and,
(v) directly impose functional restrictions on $\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]$.
Finally, to see how the unconditional variations in the exogenous regressors help to satisfy the required residual variation in endogenous regressors, suppose $X_{it,1}$ is endogenous while $X_{it,-1}$ (all other coordinates of $X_{it}$ except $X_{it,1}$) are exogenous. Suppose the conditional support of $X_{it,1}$ given $(V_{it},W_{i})$ only contains a non-zero singleton, while the unconditional support of $X_{it,-1}$ contains a small ball of positive radius. Denote $b_{1}\left(V_{it},W_{i}\right)=\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right]$. Then, I can identify the first-order moments of $\beta_{it}$ by taking partial derivative with respect to $X_{it,-1}$ on both sides of
$\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]=X_{it,1}b_{1,1}\left(V_{it},W_{i}\right)+X_{it,-1}^{\prime}b_{1,-1}\left(V_{it},W_{i}\right)$ (A.76)
to identify
$b_{1,-1}\left(V_{it},W_{i}\right)=\partial\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]/\partial X_{it,-1}.$ (A.77)
Finally, identification of $b_{1,1}\left(V_{it},W_{i}\right)$ can be obtained as
$b_{1,1}\left(V_{it},W_{i}\right)=\left(\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]-X_{it,-1}^{\prime}b_{1,-1}\left(V_{it},W_{i}\right)\right)/X_{it,1}.$
Note that there are two conditions on $X_{it,-1}$ needed for the argument to work. First, $X_{it,-1}$ needs to be exogenous, so that it does not enter the construction of $V_{it}$ and $W_{i}$. Because of this, when taking partial derivative with respect to $X_{it,-1}$ in (A.76), it directly leads to (A.77). Second, $X_{it,-1}$ needs to be continuous vector to enable the derivative-based argument. When $X_{it,-1}$ is discrete, one may use arithmetic differencing technique to identify $b_{1,-1}\left(V_{it},W_{i}\right)$.
Appendix B Additional Empirical Results
I conduct several robustness checks of the empirical findings in Section 5.
Measurement error. To address measurement error that is typical in firm-level data, I go beyond the baseline cleaning (removing outliers, negative, and extreme values) and re-estimate the model using alternative deflators and excluding firms in the top and bottom 1% of total output. Table A.1 reports the results; the estimates largely remain in line with the baseline.
Instrument validity. I probe instrument validity by varying the geographic aggregation (provincial rather than finer units), using unweighted competitor averages, and adding city fixed effects. As shown in Table A.2, the main conclusion is unchanged. The specifications with city fixed effects yield noisier estimates, likely because many firms have few within-city competitors, making the within-city demeaning numerically less stable.
Control-function specification. I assess sensitivity to the control-function specification along two dimensions. First, I consider an alternative construction of $W$ that includes averages of both the instruments and the regressors. Second, I re-estimate the series control function using alternative bases: second-degree polynomials and third-degree cubic splines with a knot at the median. Table A.3 shows that the estimates are essentially unchanged. One systematic difference is that cubic splines tend to deliver larger point estimates and wider confidence intervals across industries—consistent with the simulation evidence in Table A.8—suggesting a curse-of-dimensionality cost when the basis becomes too rich.
Comparison with classic methods. I further compare the TERC estimates with standard OLS and a naive IV approach in Table A.4. Overall, the TERC method yields higher capital elasticities and lower labor elasticities than OLS in three sectors; in the remaining two sectors, the estimates are broadly similar. These patterns echo the findings of Olley and Pakes (1996), albeit in a different context and using a different dataset. By contrast, the naive IV estimates are economically implausible, underscoring the importance of accounting for potentially nonlinear, nonseparable heterogeneity through time-varying random coefficients.
Taken together, these exercises indicate that the results are robust to a wide range of alternative specifications and sample restrictions.
Table A.1: Robustness to Measurement Error
Textile TERC Alt Deflator Non-Extreme Output
Capital Elasticity 0.415 0.403 0.408
95% CI [0.393, 0.441] [0.379, 0.428] [0.385, 0.436]
Labor Elasticity 0.452 0.472 0.441
95% CI [0.415, 0.485] [0.436, 0.505] [0.405, 0.475]
Chemical TERC Alt Deflator Non-Extreme Output
Capital Elasticity 0.463 0.420 0.463
95% CI [0.430, 0.497] [0.387, 0.453] [0.429, 0.492]
Labor Elasticity 0.311 0.439 0.293
95% CI [0.263, 0.361] [0.387, 0.489] [0.245, 0.339]
Nonmetallic Mineral TERC Alt Deflator Non-Extreme Output
Capital Elasticity 0.371 0.377 0.361
95% CI [0.332, 0.410] [0.341, 0.417] [0.324, 0.398]
Labor Elasticity 0.311 0.314 0.291
95% CI [0.255, 0.365] [0.258, 0.365] [0.238, 0.341]
General Equipment TERC Alt Deflator Non-Extreme Output
Capital Elasticity 0.433 0.434 0.422
95% CI [0.402, 0.469] [0.401, 0.474] [0.393, 0.457]
Labor Elasticity 0.521 0.526 0.522
95% CI [0.479, 0.564] [0.477, 0.569] [0.479, 0.566]
Transportation Equipment TERC Alt Deflator Non-Extreme Output
Capital Elasticity 0.425 0.453 0.429
95% CI [0.393, 0.459] [0.415, 0.493] [0.396, 0.463]
Labor Elasticity 0.588 0.566 0.571
95% CI [0.536, 0.644] [0.508, 0.632] [0.515, 0.631]
  • Notes: “Alt Deflator” refers to the price deflators constructed by Brandt et al. (2014). “Non-Extreme Output” stands for the subsample that excludes the top and bottom 1% firms in terms of total output.
Table A.2: IV Validity Check
Textile TERC Provincial IV Unweighted IV Location FE
Capital Elasticity 0.415 0.393 0.414 0.410
95% CI [0.393, 0.441] [0.368, 0.417] [0.390, 0.442] [0.386, 0.436]
Labor Elasticity 0.452 0.458 0.460 0.427
95% CI [0.415, 0.485] [0.423, 0.494] [0.425, 0.494] [0.392, 0.462]
Chemical TERC Provincial IV Unweighted IV Location FE
Capital Elasticity 0.463 0.444 0.459 0.449
95% CI [0.430, 0.497] [0.411, 0.477] [0.425, 0.492] [0.409, 0.490]
Labor Elasticity 0.311 0.321 0.303 0.360
95% CI [0.263, 0.361] [0.274, 0.371] [0.257, 0.352] [0.300, 0.411]
Nonmetallic Mineral TERC Provincial IV Unweighted IV Location FE
Capital Elasticity 0.371 0.309 0.346 0.410
95% CI [0.332, 0.410] [0.258, 0.353] [0.303, 0.387] [0.370, 0.455]
Labor Elasticity 0.311 0.289 0.321 0.381
95% CI [0.255, 0.365] [0.227, 0.345] [0.263, 0.377] [0.321, 0.429]
General Equipment TERC Provincial IV Unweighted IV Location FE
Capital Elasticity 0.433 0.433 0.434 0.414
95% CI [0.402, 0.469] [0.400, 0.467] [0.404, 0.471] [0.378, 0.452]
Labor Elasticity 0.521 0.535 0.520 0.572
95% CI [0.479, 0.564] [0.492, 0.581] [0.479, 0.563] [0.523, 0.618]
Transportation Equipment TERC Provincial IV Unweighted IV Location FE
Capital Elasticity 0.425 0.408 0.424 0.382
95% CI [0.393, 0.459] [0.372, 0.440] [0.393, 0.460] [0.346, 0.418]
Labor Elasticity 0.588 0.589 0.587 0.612
95% CI [0.536, 0.644] [0.538, 0.649] [0.538, 0.645] [0.558, 0.673]
  • Notes: “Provincial IV” denotes the leave-one-out weighted input prices of the competitors in the same province (c.f., baseline IV constructed at city level), with the weights given by the firm’s total debt (for interest rate) and number of employees (for wages). “Unweighted IV” stands for the leave-one-out unweighted input prices of the competitors in the same city. “Location FE” includes a city fixed effect in the three-step series regression.
Table A.3: Sensitivity to Control-Function Specification
Textile TERC Alt $W$ Alt Basis I Alt Basis II
Capital Elasticity 0.415 0.416 0.385 0.445
95% CI [0.393, 0.441] [0.394, 0.443] [0.364, 0.407] [0.423, 0.478]
Labor Elasticity 0.452 0.448 0.438 0.486
95% CI [0.415, 0.485] [0.413, 0.482] [0.406, 0.469] [0.451, 0.528]
Chemical TERC Alt $W$ Alt Basis I Alt Basis II
Capital Elasticity 0.463 0.462 0.434 0.490
95% CI [0.430, 0.497] [0.429, 0.495] [0.402, 0.464] [0.450, 0.534]
Labor Elasticity 0.311 0.312 0.250 0.349
95% CI [0.263, 0.361] [0.265, 0.362] [0.206, 0.291] [0.297, 0.406]
Nonmetallic Mineral TERC Alt $W$ Alt Basis I Alt Basis II
Capital Elasticity 0.371 0.357 0.334 0.370
95% CI [0.332, 0.410] [0.317, 0.396] [0.298, 0.369] [0.320, 0.415]
Labor Elasticity 0.311 0.303 0.331 0.303
95% CI [0.255, 0.365] [0.245, 0.356] [0.280, 0.386] [0.239, 0.358]
General Equipment TERC Alt $W$ Alt Basis I Alt Basis II
Capital Elasticity 0.433 0.428 0.407 0.442
95% CI [0.402, 0.469] [0.398, 0.464] [0.377, 0.441] [0.408, 0.484]
Labor Elasticity 0.521 0.513 0.496 0.530
95% CI [0.479, 0.564] [0.471, 0.556] [0.455, 0.540] [0.486, 0.579]
Transportation Equipment TERC Alt $W$ Alt Basis I Alt Basis II
Capital Elasticity 0.425 0.424 0.411 0.416
95% CI [0.393, 0.459] [0.393, 0.458] [0.383, 0.445] [0.376, 0.451]
Labor Elasticity 0.588 0.586 0.525 0.616
95% CI [0.536, 0.644] [0.539, 0.644] [0.471, 0.574] [0.563, 0.678]
  • Notes: “Alt WW” means including the time average of both the IVs and regressors into the $W$ vector. “Alt Basis I” represents the second-degree polynomial basis function. “Alt Basis II” denotes the third-degree spline basis function.
Table A.4: Comparison with OLS and Naive IV Estimates
Textile TERC OLS Naive IV
Capital Elasticity 0.415 0.379 1.309
95% CI [0.393, 0.441] [0.358, 0.400] [0.889, 1.729]
Labor Elasticity 0.452 0.509 0.266
95% CI [0.415, 0.485] [0.481, 0.537] [-0.281, 0.813]
Chemical TERC OLS Naive IV
Capital Elasticity 0.463 0.452 0.555
95% CI [0.430, 0.497] [0.419, 0.484] [-0.094, 1.204]
Labor Elasticity 0.311 0.339 -0.757
95% CI [0.263, 0.361] [0.291, 0.386] [-1.302, -0.212]
Nonmetallic Mineral TERC OLS Naive IV
Capital Elasticity 0.371 0.382 -0.670
95% CI [0.332, 0.410] [0.354, 0.411] [-4.605, 3.265]
Labor Elasticity 0.311 0.384 -3.512
95% CI [0.255, 0.365] [0.340, 0.427] [-9.599, 2.575]
General Equipment TERC OLS Naive IV
Capital Elasticity 0.433 0.389 0.951
95% CI [0.402, 0.469] [0.363, 0.416] [0.765, 1.138]
Labor Elasticity 0.521 0.501 -0.619
95% CI [0.479, 0.564] [0.464, 0.538] [-1.220, -0.019]
Transportation Equipment TERC OLS Naive IV
Capital Elasticity 0.425 0.442 0.723
95% CI [0.393, 0.459] [0.408, 0.475] [0.523, 0.924]
Labor Elasticity 0.588 0.560 0.508
95% CI [0.536, 0.644] [0.509, 0.612] [-0.110, 1.127]
  • Notes: "OLS" denotes the pooled OLS method with constant coefficients to estimate the output elasticities. "Naive IV" represents two-stage least squares method where the IVs are the same as in TERC estimates. CIs are obtained using standard Stata command.
@@ Appendix C Simulation
In this section, I examine the finite-sample performance of the series estimators via a simulation study. A discussion of the data generating process (DGP) motivated by production function applications is first provided. Then, I present the baseline results of estimating the APE $\overline{b}\coloneqq T^{-1}\sum_{t=1}^{T}\mathbb{E}\beta_{it}$. Next, I show how the method performs for estimating the LAR $b_{t}\left(x\right)\coloneqq\mathbb{E}\left[\left.\beta_{it}\right|X_{it}=x\right]$. Finally, as robustness checks, I (i) vary the number of firms and periods, (ii) change the degree of basis functions used for series estimation, and (iii) include ex-post shocks $\widetilde{\epsilon}_{it}$ and $\upsilon_{it}$.
@@ C.1 DGP
$Y_{it}=k_{it}\beta_{it,K}+l_{it}\beta_{it,L}+\omega_{it}+\widetilde{\epsilon}_{it},$
where $\omega_{it},\ \beta_{it,K},\ \text{and }\beta_{it,L}$ are all functions of $\left(A_{i},\varepsilon_{it},\upsilon_{it}\right)$, $k_{it}$ and $l_{it}$ are the natural logs of optimal capital and labor calculated from the solution to firm $i$’s profit maximization problem, and $Y_{it}$ is the natural log of value-added output measured in dollars. Let $\mathcal{U}\left[a,b\right]$ denote the uniform distribution over $\left[a,b\right]$. I draw $A_{i}\sim_{i.i.d.}\mathcal{U}\left[1,2\right]$ and $\varepsilon_{it}\sim_{i.i.d.}\mathcal{U}\left[1,2\right]$. I let true $\omega_{it}=\ln\left(A_{i}+\varepsilon_{it}/2+1\right)$, $\beta_{it,K}=\left(A_{i}+\varepsilon_{it}\right)/10$, and $\beta_{it,L}=\left(A_{i}+\varepsilon_{it}\right)/10$, and write $\beta_{it}=\left(\omega_{it},\beta_{it,K},\beta_{it,L}\right)^{\prime}$. I compute the true $\overline{\omega}\coloneqq T^{-1}\sum_{t=1}^{T}\mathbb{E}\omega_{it}=1.1736,$ $\overline{b}_{K}\coloneqq T^{-1}\sum_{t=1}^{T}\mathbb{E}\beta_{it,K}=.3$, and $\overline{b}_{L}\coloneqq T^{-1}\sum_{t=1}^{T}\mathbb{E}\beta_{it,L}=.3$. The range of the random coefficients are set such that the second-order condition for the profit maximization problem is satisfied. For the baseline results, I let ex-post shocks $\widetilde{\epsilon}_{it}=\upsilon_{it}=0$ and investigate their impacts as robustness checks later. Suppose capital and labor choices are made separately (e.g., investment and hiring decisions made by different departments) based on ex-ante signals about $\varepsilon_{it}$ with noise $\lambda_{it,K}$ and $\lambda_{it,L}$: $\eta_{it,K}=\varepsilon_{it}+\lambda_{it,K}$ and $\eta_{it,L}=\varepsilon_{it}+\lambda_{it,L}$, where $\lambda_{it,K}\text{ and }\lambda_{it,L}\sim_{i.i.d.}\mathcal{U}\left[-.05,.05\right]$. Since $Y_{it}$ is measured in dollars which is the case with most real production datasets, the price of output $P_{it}$ is assumed to be 1. I draw each element of the IVs $Z_{it}=\left(r_{it},w_{it}\right)^{\prime}$ from $\mathcal{U}\left[0,\ln 3\right]$ independent of each other and all other variables (in particular, $Z_{it}\perp(\varepsilon_{it},\eta_{it})^{\prime}$), and solve the firm’s profit maximization problem to obtain
$k_{it}$ $=\dfrac{r_{it}-\left(A_{i}+\eta_{it,K}\right)\left(r_{it}-w_{it}\right)/10-\ln\left(\left(A_{i}+\eta_{it,K}\right)/10\right)-\ln\left(A_{i}+\eta_{it,K}/2+1\right)}{\left(A_{i}+\eta_{it,K}\right)/5-1},$
$l_{it}$ $=\dfrac{w_{it}-\left(A_{i}+\eta_{it,L}\right)\left(w_{it}-r_{it}\right)/10-\ln\left(\left(A_{i}+\eta_{it,L}\right)/10\right)-\ln\left(A_{i}+\eta_{it,L}/2+1\right)}{\left(A_{i}+\eta_{it,L}\right)/5-1}.$
Let $X_{it}=\left(k_{it},l_{it}\right)^{\prime}$ denote the two endogenous choice variables. It is clear that $X_{it}$ is correlated with $\beta_{it}$ in each period via $\left(A_{i},\eta_{it,K},\eta_{it,L}\right)$ in a nonseparable way. I use $n$, $T$, and $M$ to denote the total number of firms, periods, and simulations, respectively. The simulated dataset consists of $\left\{X_{it}^{(m)},Y_{it}^{(m)},Z_{it}^{(m)}\right\}$ for $i=1,\ldots,n$, $t=1,\ldots,T$, and $m=1,\ldots,M$, which are used to construct $\widehat{\overline{b}}^{(m)}$ and $\widehat{b}_{t}^{(m)}\left(x\right)$ to estimate $\overline{b}$ and $b_{t}\left(x\right)$, respectively, via the estimation procedure outlined in Subsection 4.1. I evaluate the performance of the estimators by their biases, root-mean squared errors (rMSE), and mean normed deviations (MND), with the explicit mathematical definitions provided in the tables below.